OpenStudy (anonymous):

The space shuttle travels at 28,000 mph wile orbiting the earth how far does the shuttle travel in feet in a day

OpenStudy (anonymous):

thats not you in the picture, right?..

OpenStudy (anonymous):

\[28000\frac{ mi }{ h }* \frac{ 24h }{ d }* \frac{ 5280feet }{ mi } = x \frac{ feet }{ day }\]

OpenStudy (anonymous):

There are 24 hours in a day, and if it is travelling 28,000 miles per hours, then it would be 24*28000.

OpenStudy (anonymous):

hour*

OpenStudy (anonymous):

3.55*10^9

OpenStudy (anonymous):

Then since there is 5,280 feet in a mile, you would divide that by 5,280.